OpenStudy (wavykelp):

If f(x) varies directly with x2, and f(x) = 75 when x = 5, find the value of f(4). A. 100 B. 50 C. 48 D. 25

OpenStudy (wavykelp):

is it A?

jimthompson5910 (jim_thompson5910):

" f(x) varies directly with x^2", so f(x) = k*x^2 for some constant k

jimthompson5910 (jim_thompson5910):

we know that "f(x) = 75 when x = 5", so... f(x) = k*x^2 75 = k*(5)^2 75 = k*25 75/25 = k 3 = k k = 3

OpenStudy (wavykelp):

ohhhhhh i get it

jimthompson5910 (jim_thompson5910):

Which makes the equation turn from f(x) = k*x^2 into f(x) = 3x^2

jimthompson5910 (jim_thompson5910):

now that you know the full function, you can plug in x = 4 and evaluate

OpenStudy (wavykelp):

so its 100....

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

plug x = 4 into f(x) = 3x^2 to get ???

OpenStudy (wavykelp):

144

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

you square before you multiply

jimthompson5910 (jim_thompson5910):

exponents come before multiplication in PEMDAS

OpenStudy (wavykelp):

ohh of course

OpenStudy (wavykelp):

its 48

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (wavykelp):

thx

jimthompson5910 (jim_thompson5910):

you're welcome